# 25mm25mm25mm26mm

### v2

@@ 4. Zig flow
Recall that in the bulk regime $\gamma_{t}\sim|z_{1,t}-z_{2,t}|^{2}+|\eta_{1,t}|+|\eta_{2,t}|$. From the monotonicity of $t\mapsto|\eta_{t}|$ it also follows that $\gamma_{s}\gtrsim\gamma_{t}$ for $s\leq t$. If $|z_{1,t}-z_{2,t}|\geq n^{-\delta}$, for some very small fixed $100<\delta\leq\xi/$ 10, then the desired result follows by a simple (almost) global law (see e.g. [28, Lemma 5.4]). For the remainder of the proof we thus assume $|z_{1,t}-z_{2,t}|\leq n^{-\delta}$. Note that $|z_{1,t}-z_{2,t}|\sim|z_{1,s}-z_{2,s}|$ for any $s,t$, by (4.9), it thus makes no difference at which time we impose conditions on $|z_{1,t}-z_{2,t}|$.
@@
and proceeding similarly to [31, Eqs. (5.28)–(5.41)], we readily obtain
@@
Here $e_{t},h_{t}$ are a martingale and a forcing term, respectively, such that
@@
with very high probability. Here $[\cdot]_{t}$ denotes the quadratic variation process.
By explicit, but fairly tedious, computations we find (see (A) in Appendix A) (recall $\lVert A_{i}\rVert\leq 1$)
@@
From (4.25), we thus obtain
@@
Finally, by (4.29), using a Gronwall inequality together with (4.18) to estimate $Y_{0}\leq\alpha_{t\wedge\tau_{1}}$, we obtain $Y_{t\wedge\tau_{1}}\leq n^{\xi}\alpha_{t\wedge\tau_{1}}$. We point out that here we also used
@@
For products of three resolvents we have:
@@
Here $\widetilde{\gamma}_{t}:=\gamma(\widetilde{\eta}_{1,t},z_{1,t},\widetilde{\eta}_{2,t},z_{2,t})$, with $\widetilde{\eta}_{i,t}$ being the solution of (4.5) with initial condition $\widetilde{\eta}_{i,0}$, and ~$\widetilde{\ell}_{t}$, $\widetilde{\eta}_{*,t}$ are defined similarly. The supremum $\sup^{*}$ is taken over all the initial conditions $\widetilde{\eta}_{i,0}$ with $\widetilde{\eta}_{i,0}\in[\eta_{i,0},\omega_{*}]$. We point out that in (4.40) we need to take the supremum over a larger set of initial conditions $\widetilde{\eta}_{i,0}$ only because at some point (see (4.60) below) we need to consider quantities involving $|G|$, and for $|G|$ we use an integral representation which requires integration over an entire segment of spectral parameters.
@@
with very high probability. Here $\check{\gamma}_{t}:=\gamma(\check{\eta}_{1},z_{1,t},\check{\eta}_{2},z_{2,t})$, and ˇ$\check{\ell}_{t}$, $\check{\eta}_{*,t}$ are defined similarly. We point that in (4.41) we also used that, as a consequence of (4.2), for any fixed $t\leq T_{1}^{\epsilon}\wedge T_{2}^{\epsilon}$ and any $\check{\eta}_{i}\in[\eta_{i,0},\omega_{*}/2]$ there is a $\widetilde{\eta}_{i,0}\in[\eta_{i,0},\omega_{*}]$ such that $\widetilde{\eta}_{i,t}=\check{\eta}_{i}$.
@@
By the Burkholder–Davis–Gundy inequality we thus obtain
@@
Estimates for (LABEL:eq:3gisosubma): Similarly to the two resolvent case above, we notice that the second term in the rhs. of (LABEL:eq:3gisosubma) can be neglected from the analysis, since it only amounts to a simple rescaling of size $e^{3t/2}\sim 1$. Several of the following estimates are similar to the estimates that we preformed in the rhs. of (4.31), for this reason we omit some easily checkable details.
@@
where $\widehat{\mathcal{C}}_{t}$ denotes the $4\times 4$ covariance matrix of the martingale ^$\widehat{\mathcal{E}}_{t}$. Here we also used the notations
@@
and $b_{t}$ is real and defined in [31, Eq. (A.14)]. Note that the matrix ^$\widehat{\mathcal{M}}_{t}$ is exactly the same as the one [31, Eq. (5.44)], and that $\alpha_{t\wedge\tau_{2}}$ defined in (4.68) is different compared to the one defined in (4.54). We can thus use again the matrix Gronwall inequality from [31, Lemma 5.6] for $\alpha_{t}$ defined in (4.68) and $f_{t}:=2\langle M_{12,t}^{E_{+}}\rangle$ (note that here we redefined $f_{t}$, i.e. there is an additional factor of $2$ compared to (4.54)). By (4.55), we thus conclude
@@ 4.1. Improved local law at the edge of the spectrum
In this section we present the proof of the averaged local law at the edge of the spectrum, including $\mathrm{Im}\,G$’s. While in the bulk of the spectrum there is no difference between $G$ and $\mathrm{Im}\,G$, we will see that close to the edge there will be an improvement (proportional to the local density) if we replace one (or more) $G$ by $\mathrm{Im}\,G$. For brevity, we present this proof only in the averaged case, since this is what gives the optimal bound on the eigenvector overlaps in Theorem 3.6.
@@
Fix small $n$–independent constants $\tau,\omega_{*},\epsilon>0$, and for $i=1,2$ fix spectral parameters $\Lambda_{i,0}$ as in (4.12), with $|z_{i,0}|\leq 10$, $|\eta_{i,0}|\leq\omega_{*}$. For $t\leq T^{*}$ let $\Lambda_{i,t}$ be the solution of (4.4) with initial condition $\Lambda_{i,0}$. Let $G_{i,t}:=(W_{t}-\Lambda_{i,t})^{-1}$, and let $\widehat{M}_{12,t}^{A}$ be the deterministic approximation of $\mathrm{Im}\,G_{1,t}A\mathrm{Im}\,G_{2,t}$ from (3.27). Let $\ell$ from (4.13), and define the control parameter
@@
Assume that for some small $100<\xi\leq\epsilon/$ 10, with very high probability, it holds
@@
5. The zag step: Proof of Theorem 3.4
@@
Fix small constants $\epsilon,\tau>0$. Assume that for any $z_{i}\in{\mathbb{C}}$ and $\eta_{i}\in{\mathbb{R}}$ satisfying $|z_{i}|\leq 1+\tau$, $\min_{i=1}^{2}\{|\eta_{i}|\rho_{i}\}\geq n^{-1+\epsilon}$, $\gamma\gtrsim\eta_{*}/\rho^{*}$, the following hold for some $10\leq b<$ 1, i.e.,
@@
uniformly for any deterministic bounded matrices $A_{i}~{}(i=1,2)$ and unit vectors ${\bm{x}},{\bm{y}}$, where $\gamma$ is defined in (3.23) and $M_{12}^{A_{1}}$, $M_{121}^{A_{1},A_{2}}$ are defined in (3.16), (3.19). Then the following also hold for a larger $b^{\prime}=\min\big{\{}b+\theta,~{}1\big{\}}$ with some $100<\theta\leq\epsilon/$ 10,
@@
assuming that $|R^{xy}|\prec\mathcal{E}_{b}$ and $|S^{xy}|\prec\widetilde{\mathcal{E}}_{b}$ for some $10\leq b<$ 1.
The proof of Proposition 5.2 will be a Gronwall argument [zag step] together with Proposition 4.4 [zig step]. The bounds (5.6) and (5.7) are already known from Proposition 4.4 (together with the global laws needed as an initial input that will be stated later in (5.18)–(5.19)) whenever $X$ has a small Gaussian component. We show that this Gaussian component does not change the estimate up to the relevant order by running an Ornstein-Uhlenbeck flow to remove the small Gaussian component. We will work with the time evolution of high moments of $R^{xy}_{t},S^{xy}_{t}$. The key of the Gronwall argument is to give an estimate on the time derivative of these high moments in terms of themselves, this will be the content of Lemma 5.4 below.
@@
The corresponding resolvent is denoted by $G^{z}_{t}(\mathrm{i}\eta):=(H_{t}^{z}-\mathrm{i}\eta)^{-1}$, which we use to define time dependent $R^{xy}_{t}$, $S^{xy}_{t}$ as in (5.9). Note that the parameters $z_{i}\in{\mathbb{C}}$, $\eta_{i}\in{\mathbb{R}}$ are fixed and hence the deterministic matrices $M^{A}_{12}$ and $M^{A_{1},A_{2}}_{121}$ in (5.9) are independent of the time $t$.
@@
Fix small constant $\epsilon,\tau>0$ and any integer $m\geq 1$. For any $|z_{i}|\leq 1+\tau$ and $\eta_{i}\in{\mathbb{R}}$ satisfying $\min_{i=1}^{2}\{|\eta_{i}|\rho_{i}\}\geq n^{-1+\epsilon}$ and $\gamma\gtrsim\eta_{*}/\rho^{*}$, assuming [^3] We remark that stochastic domination $\prec$ and upper bounds of arbitrarily high moments of a random variable are equivalent for random variables bounded by some $n^{C}$ which is always true in our case, i.e., for $|X|\leq n^{C}$ we have that $X\prec\Psi$ if and only if ${\mathbf{E}}|X|^{2m}\leq n^{\xi}\Psi^{2m}$ for any $m\geq 1$ and $\xi>0$. We frequently use this fact in the proof without mentioning it. that for some $10\leq b<$ 1,
@@
uniformly for any $t\geq 0$, any deterministic bounded matrices $A_{i}~{}(i=1,2)$ and unit vectors ${\bm{x}},{\bm{y}}$, then the following hold for $b^{\prime}=\min\big{\{}b+\theta,~{}1\big{\}}$, $100<\theta\leq\epsilon/$ 10, and the prefactor $C_{n}:=1+\frac{1}{\sqrt{n}\eta_{*}^{1+\frac{\epsilon}{2}}}$,
@@
To extend the above results to smaller $|\eta_{*}|\lesssim n^{-1/2+\epsilon/2}$, we use zig-zag iteratively as follows (see the illustration in Figure 1 below). In each iteration step $k\geq 1$, we use the superscript $(k)$ to indicate the step number, e.g., we denote the spectral parameters by $\eta^{(k)}_{i}\in{\mathbb{R}}$, $z^{(k)}_{i}\in{\mathbb{C}}$ and the time by $T^{(k)}\in{\mathbb{R}}^{+}$. Recall that our final goal is to prove (5.12) for any fixed $|z_{i}|\leq 1+\tau$ and $n|\eta_{i}|\rho_{i}\gtrsim n^{\epsilon}$. We also recall from Lemma 4.1 that the size of $\rho_{i}$ does not change along the characteristic flow in (4.4) by more than a constant factor. Note that (3.10) and $n|\eta_{i}|\rho\gtrsim n^{\epsilon}$ imply that $|\eta_{i}|\gtrsim n^{-1+\epsilon}$ and $n^{-1/2+\epsilon/2}\lesssim\rho_{i}\lesssim 1$. Then we set $\rho_{*}=\min_{i=1}^{2}\rho_{i}\sim n^{-c_{*}}$ with $20\leq c_{*}\leq 1/2-\epsilon/$ 2. In general for any $k\geq 1$, we define the scales recursively as follows:
@@
We will stop the iterations if $\eta^{(k)}_{*}$ reaches $\eta_{*}=\min_{i=1}^{2}|\eta_{i}|\gtrsim n^{-1+\epsilon}$, so $a_{k}\geq 1-\epsilon$. The number of iteration steps is denoted by $K$. It is straightforward to check that $a_{k}$ is increasing in $k$ at least up to $1$ using that $20\leq c_{*}\leq 1/2-\epsilon/$ 2.
In the $k$-th step, we fix $|\eta_{i}^{(k)}|\gtrsim\eta_{*}^{(k)}=n^{-a_{k}}$ and $|z^{(k)}_{i}|\leq 1+\tau$ with $\rho_{i}^{(k)}\sim\rho_{i}$. Assume that in the previous step, we have proved (5.22) for any $|\eta^{(k-1)}_{i}|\gtrsim\eta_{*}^{(k-1)}=n^{-a_{k-1}}$ and $|z^{(k-1)}_{i}|\leq 1+\tau$ with $\rho_{i}^{(k-1)}\sim\rho_{i}$. Recall the characteristic flow $\Lambda(t)=\Lambda_{t}$ in (4.4) and choose $T^{(k)}\sim\eta_{*}^{(k-1)}/\rho_{*}=n^{-a_{k-1}+c_{*}}$. From Lemma 4.1, there exist initial conditions $|z_{i}(0)|=|z_{i}^{(k)}|+O(T^{(k)})$ and $|\eta_{i}(0)|\gtrsim\rho_{*}T^{(k)}=\eta_{*}^{(k-1)}=n^{-a_{k-1}}$ such that $\eta_{i}(T^{(k)})=\eta^{(k)}_{i}$ and $z_{i}(T^{(k)})=z^{(k)}_{i}$. The result from the previous step in (5.22) ensures that the assumption (4.18) of Proposition 4.4 (with $b\textquoteright$ ’ playing the role of $b$) is satisfied for any $|\eta^{(k-1)}_{i}|\gtrsim\eta_{*}^{(k-1)}=n^{-a_{k-1}}$. Thus Proposition 4.4 implies that for $t=T^{(k)}\sim n^{-a_{k-1}+c_{*}}$,
@@
We first apply Itô’s formula to $|R^{xy}_{t}|^{2m}=(R^{xy}_{t})^{m}((R^{xy}_{t})^{*})^{m}$ and perform the cumulant expansions (see e.g., [56, Lemma 7.1]) as below. Without loss of generality, we may assume in the following that all the cumulants of the normalized entries of $\sqrt{n}X$ are equal to one. We thus obtain
@@
$\mathcal{T}^{(s_{1},\cdots,s_{l})}_{k}:=\frac{1}{n^{\frac{k}{2}}}\sum_{a=1}^{n}\sum_{B=n+1}^{2n}{\mathbf{E}}\Big{[}\big{(}\partial^{s_{1}}R^{xy}_{t}\big{)}\cdots\big{(}\partial^{s_{l}}R^{xy}_{t}\big{)}\big{(}R^{xy}_{t}\big{)}^{2m-l}\Big{]},\qquad s_{i}\geq 1.$ (5.27)
Here we stopped the cumulant expansions in (5.1) at a sufficiently high power $K_{0}=O(m)$, and the last error term $O_{\prec}(n^{-100m})$ was estimated using the moment condition in (2.1) and that $\|G\|\leq|\eta|^{-1}$. This argument is quite standard and already used in many previous works, e.g., below [31, Eq. (7.6)], so we omit the details. For notational brevity, we abuse the usage of the notation $R^{xy}_{t}$ to indicate either $R^{xy}_{t}$ itself or its complex conjugate $(R^{xy}_{t})^{*}$. We also use the notation $\partial$ to indicate either $\partial/\partial h_{aB}$ or $\partial/\partial h_{Ba}$.
@@
with $b^{\prime}=\min\big{\{}b+\theta,~{}1\big{\}}$, $100<\theta\leq\epsilon/$ 10, using that $\gamma\lesssim\rho^{-2}$ from (3.23) and $\gamma\gtrsim\eta/\rho$.
@@
|()|(*)$\prec$| ( ∗ ) | ≺ $\frac{\rho^{3/2}}{n\sqrt{\eta}}{\mathbf{E}}\Big{[}\sqrt{\sum_{B}\big{|}(G_{1})_{{\bm{x}}{\bm{e}}_{B}}\big{|}^{2}\sum_{B}\big{|}(G_{1}AG_{2})_{{\bm{e}}_{B}{\bm{y}}}\big{|}^{4}}\big{|}R^{xy}_{t}\big{|}^{2m-2}\Big{]}$

### v3

@@ 4. Zig flow
Recall that by (4.17) in the bulk regime $\gamma_{t}\sim|z_{1,t}-z_{2,t}|^{2}+|\eta_{1,t}|+|\eta_{2,t}|$. From the monotonicity of $t\mapsto|\eta_{t}|$ (see Lemma 4.1), and (4.9) it also follows that $\gamma_{s}\gtrsim\gamma_{t}$ for $s\leq t$. If $|z_{1,t}-z_{2,t}|\geq n^{-\delta}$, for some very small fixed $0<\delta\leq\xi/10$, then the desired result follows by a simple (almost) global law (see e.g. [28, Lemma 5.4]). For the remainder of the proof we thus assume $|z_{1,t}-z_{2,t}|\leq n^{-\delta}$. Note that $|z_{1,t}-z_{2,t}|\sim|z_{1,s}-z_{2,s}|$ for any $s,t$, by (4.9), it thus makes no difference at which time we impose conditions on $|z_{1,t}-z_{2,t}|$.
@@
and proceeding similarly to [31, Eqs. (5.28)–(5.41)], from (4.20) we readily obtain
@@
Here $e_{t},h_{t}$ are a martingale and a forcing term such that
@@
with very high probability. Here $\cdot]_{t}$ denotes the quadratic variation process.
By explicit but fairly tedious computations we find (see (A) in Appendix A) (recall $\lVert A_{i}\rVert\leq 1$)
@@
Integrating (4.25) in time (from 0 to $t\wedge\tau_{1}$) and using (4.27) to estimate the $M_{12,t},M_{21,t}$–terms in (4.25), we thus obtain
@@
Finally, by (4.29), using a Gronwall inequality together with (4.18) to estimate $Y_{0}\leq\alpha_{t\wedge\tau_{1}}$, we obtain $Y_{t\wedge\tau_{1}}\leq n^{\xi}\alpha_{t\wedge\tau_{1}}$. We point out that here we also used (recall the definition of $\alpha_{t}$ from (4.23))
@@
Here we recall that $\partial_{ab}$ denotes the directional derivative in the direction $(W_{t})_{ab}$, $B_{ab,t}$ denotes the (a,b)(a,b)–entry of the matrix Brownian motion $B_{t}$ from (4.1), and $\sum_{i\neq j}$ denotes a summation over the pairs $(i,j)\in\{(1,2),(2,1)\}$. For products of three resolvents we have:
@@
Here $\widetilde{\gamma}_{t}:=\gamma(\widetilde{\eta}_{1,t},z_{1,t},\widetilde{\eta}_{2,t},z_{2,t})$, with $\widetilde{\eta}_{i,t}$ being the solution of (4.5) with initial condition $\widetilde{\eta}_{i,0}$, and $\widetilde{\ell}_{t}$, $\widetilde{\eta}_{*,t}$ are defined similarly. The supremum $\sup^{*}$ is taken over all the initial conditions $\widetilde{\eta}_{i,0}$ with $\widetilde{\eta}_{i,0}\in[\eta_{i,0},\omega_{*}]$. We point out that in (4.40) we need to take the supremum over a larger set of initial conditions $\widetilde{\eta}_{i,0}$ only because at some point (see (4.60) below) we need to consider quantities involving $|G|$, and for $|G|$ we use an integral representation which requires integration over an entire segment of spectral parameters.
@@
with very high probability. Here $\check{\gamma}_{t}:=\gamma(\check{\eta}_{1},z_{1,t},\check{\eta}_{2},z_{2,t})$, and $\check{\ell}_{t}$, $\check{\eta}_{*,t}$ are defined similarly. We point that in (4.41) we also used that, as a consequence of (4.2), for any fixed $t\leq T_{1}^{\epsilon}\wedge T_{2}^{\epsilon}$ and any $\check{\eta}_{i}\in[\eta_{i,0},\omega_{*}/2]$ there is a $\widetilde{\eta}_{i,0}\in[\eta_{i,0},\omega_{*}]$ such that $\widetilde{\eta}_{i,t}=\check{\eta}_{i}$.
@@
We point out that in the following we will often use several properties of the characteristics from Lemma 4.1 even if not stated explicitly. By the Burkholder–Davis–Gundy inequality we thus obtain
@@
Estimates for (LABEL:eq:3gisosubma): Similarly to the two resolvent case above, we notice that the second term in the rhs. of (LABEL:eq:3gisosubma) can be neglected from the analysis, since it only amounts to a simple rescaling of size $e^{3t/2}\sim 1$. Several of the following estimates are similar to the estimates that we preformed in the rhs. of (4.31), for this reason we omit some easily checkable details. In particular, we point out that throughout we will use properties of the characteristics from Lemma 4.1 even if not stated explicitly.
@@
where $\widehat{\mathcal{C}}_{t}$ denotes the $4\times 4$ covariance matrix of the martingale $\widehat{\mathcal{E}}_{t}$. Here we also used the notations
@@
and $b_{t}$ is real and defined in [31, Eq. (A.14)]. Note that the matrix $\widehat{\mathcal{M}}_{t}$ is exactly the same as the one [31, Eq. (5.44)], and that $\alpha_{t\wedge\tau_{2}}$ defined in (4.68) is different compared to the one defined in (4.54). We can thus use again the matrix Gronwall inequality from [31, Lemma 5.6] for $\alpha_{t}$ defined in (4.68) and $f_{t}:=2\langle M_{12,t}^{E_{+}}\rangle$ (note that here we redefined $f_{t}$, i.e. there is an additional factor of $2$ compared to (4.54)). By (4.55), we thus conclude
@@ 4.1. Improved local law at the edge of the spectrum
In this section we present the proof of the averaged local law at the edge of the spectrum, including $\mathrm{Im}\,G$’s. While in the bulk of the spectrum there is no difference between $G$ and $\mathrm{Im}\,G$, we will see that close to the edge there will be an improvement (proportional to the local density) if we replace one (or more) $G$ by $\mathrm{Im}\,G$. For brevity, we present this proof only in the averaged case, since this is what gives the optimal bound on the eigenvector overlaps in Theorem 3.6. The result from Proposition 4.8 below is the main technical input to prove Theorem 3.5 (see the beginning of Section 4 for a related explanation).
@@
Fix small nn–independent constants $\tau,\omega_{*},\epsilon>0$, and for $i=1,2$ fix spectral parameters $\Lambda_{i,0}$ as in (4.12), with $|z_{i,0}|\leq 10$, $|\eta_{i,0}|\leq\omega_{*}$. For $t\leq T^{*}$ let $\Lambda_{i,t}$ be the solution of (4.4) with initial condition $\Lambda_{i,0}$. Let $G_{i,t}:=(W_{t}-\Lambda_{i,t})^{-1}$, and let $\widehat{M}_{12,t}^{A}$ be the deterministic approximation of $\mathrm{Im}\,G_{1,t}A\mathrm{Im}\,G_{2,t}$ from (3.27). Let $\ell$ from (4.13), and define the control parameter
@@
Assume that for some small $0<\xi\leq\epsilon/10$, with very high probability, it holds
@@
5. The zag step: Proofs of Theorem 3.4 and Theorem 3.5
In Section 4 (zig step), we proved in Proposition 4.3 that if the local laws from Theorem 3.4 hold on a global scale $\eta\sim 1$, then they also hold down to the optimal local scale $\eta\gg 1/n$, at the cost of adding a Gaussian component. This section (zag step) is devoted to remove this added Gaussian component using a Gronwall argument (see Lemma 5.4 and (5.49) below) and hence prove Theorem 3.4. The proof of Theorem 3.5 near the edge is similar to Theorem 3.4 using (5.75) below to remove the Gaussian component added in Proposition 4.8.
The outline of this section is as follows. We first present the proof of the isotropic local law (3.25) and the averaged law in Section 5.1 and 5.2 respectively. In Sections 5.3, we prove standard local laws without $|z_{1}-z_{2}|$ decay stated in Proposition 5.1 below, which serve as an input to prove Theorem 3.4. The proof of Theorem 3.5 is presented in Section 5.4 using Theorem 3.5 as an input.
@@
Fix small constants $\epsilon,\tau>0$. Assume that for any $z_{i}\in{\mathbb{C}}$ and $\eta_{i}\in{\mathbb{R}}$ satisfying $|z_{i}|\leq 1+\tau$, $\min_{i=1}^{2}\{|\eta_{i}|\rho_{i}\}\geq n^{-1+\epsilon}$, $\gamma\gtrsim\eta_{*}/\rho^{*}$, the following hold for some $0\leq b<1$, i.e.,
@@
uniformly for any deterministic bounded matrices $A_{i}~(i=1,2)$ and unit vectors ${\bm{x}},{\bm{y}}$, where $\gamma$ is defined in (3.23) and $M_{12}^{A_{1}}$, $M_{121}^{A_{1},A_{2}}$ are defined in (3.16), (3.19). Then the following also hold for a larger $b^{\prime}=\min\big\{b+\theta,~1\big\}$ with some $0<\theta\leq\epsilon/10$,
@@
assuming that $|R^{xy}|\prec\mathcal{E}_{b}$ and $|S^{xy}|\prec\widetilde{\mathcal{E}}_{b}$ for some $0\leq b<1$.
The proof of Proposition 5.2 will be a Gronwall argument [zag step] together with Proposition 4.4 [zig step]. The bounds (5.6) and (5.7) are already known from Proposition 4.4 (together with the global laws needed as an initial input that will be stated later in (5.18)–(5.19)) whenever $X$ has a small Gaussian component. We show that this Gaussian component does not change the estimate up to the relevant order by running an Ornstein-Uhlenbeck flow to remove the small Gaussian component. We will work with the time evolution of high moments of $R^{xy}_{t},S^{xy}_{t}$. The key of the Gronwall argument is to give an estimate on the time derivative of these high moments in terms of themselves, this will be the content of Lemma 5.4 below. Similarly to the previous section, to keep the presentation simple, we now focus only on the complex case.
@@
The corresponding time-dependent resolvent of $H_{t}^{z_{i}}~(i=1,2)$ is denoted by
$G_{i}:=G^{z_{i}}_{t}(\mathrm{i}\eta_{i})=(H^{z_{i}}_{t}-\mathrm{i}\eta_{i})^{-1},\qquad i=1,2,$
which we use to define time dependent $R^{xy}_{t}$, $S^{xy}_{t}$ as in (5.9). For notational brevity, we will not keep track of the dependence of $t$ in the notations $G_{i}$ from now on. Note that the parameters $z_{i}\in{\mathbb{C}}$, $\eta_{i}\in{\mathbb{R}}$ are fixed and hence the deterministic matrices $M^{A}_{12}$ and $M^{A_{1},A_{2}}_{121}$ in (5.9) are independent of the time $t$.
@@
Fix small constant $\epsilon,\tau>0$ and any integer $m\geq 1$. For any $|z_{i}|\leq 1+\tau$ and $\eta_{i}\in{\mathbb{R}}$ satisfying $\min_{i=1}^{2}\{|\eta_{i}|\rho_{i}\}\geq n^{-1+\epsilon}$ and $\gamma\gtrsim\eta_{*}/\rho^{*}$, assuming [^5] We remark that stochastic domination $\prec$ and upper bounds of arbitrarily high moments of a random variable are equivalent for random variables bounded by some $n^{C}$ which is always true in our case, i.e., for $|X|\leq n^{C}$ we have that $X\prec\Psi$ if and only if ${\mathbf{E}}|X|^{2m}\leq n^{\xi}\Psi^{2m}$ for any $m\geq 1$ and $\xi>0$. We frequently use this fact in the proof without mentioning it. that for some $0\leq b<1$,
@@
uniformly for any $t\geq 0$, any deterministic bounded matrices $A_{i}~(i=1,2)$ and unit vectors ${\bm{x}},{\bm{y}}$, then the following hold for $b^{\prime}=\min\big\{b+\theta,~1\big\}$, $0<\theta\leq\epsilon/10$, and the prefactor $C_{n}:=1+\frac{1}{\sqrt{n}\eta_{*}^{1+\frac{\epsilon}{2}}}$,
@@
To extend the above results to smaller $|\eta_{*}|\lesssim n^{-1/2+\epsilon/2}$, we use zig-zag iteratively as follows (see the illustration in Figure 1 below). In each iteration step $k\geq 1$, we use the superscript $(k)$ to indicate the step number, e.g., we denote the spectral parameters by $\eta^{(k)}_{i}\in{\mathbb{R}}$, $z^{(k)}_{i}\in{\mathbb{C}}$ and the time by $T^{(k)}\in{\mathbb{R}}^{+}$. Recall that our final goal is to prove (5.12) for any fixed $|z_{i}|\leq 1+\tau$ and $n|\eta_{i}|\rho_{i}\gtrsim n^{\epsilon}$. We also recall from Lemma 4.1 that the size of $\rho_{i}$ does not change along the characteristic flow in (4.4) by more than a constant factor. Note that (3.10) and $n|\eta_{i}|\rho\gtrsim n^{\epsilon}$ imply that $|\eta_{i}|\gtrsim n^{-1+\epsilon}$ and $n^{-1/2+\epsilon/2}\lesssim\rho_{i}\lesssim 1$. Then we set $\rho_{*}=\min_{i=1}^{2}\rho_{i}\sim n^{-c_{*}}$ with $0\leq c_{*}\leq 1/2-\epsilon/2$. In general for any $k\geq 1$, we define the scales recursively as follows:
@@
We will stop the iterations if $\eta^{(k)}_{*}$ reaches $\eta_{*}=\min_{i=1}^{2}|\eta_{i}|\gtrsim n^{-1+\epsilon}$, so $a_{k}\geq 1-\epsilon$. The number of iteration steps is denoted by $K$. It is straightforward to check that $a_{k}$ is increasing in $k$ at least up to $1$ using that $0\leq c_{*}\leq 1/2-\epsilon/2$.
In the $k$-th step, we fix $|\eta_{i}^{(k)}|\gtrsim\eta_{*}^{(k)}=n^{-a_{k}}$ and $|z^{(k)}_{i}|\leq 1+\tau$ with $\rho_{i}^{(k)}\sim\rho_{i}$. Assume that in the previous step, we have proved (5.22) for any $|\eta^{(k-1)}_{i}|\gtrsim\eta_{*}^{(k-1)}=n^{-a_{k-1}}$ and $|z^{(k-1)}_{i}|\leq 1+\tau$ with $\rho_{i}^{(k-1)}\sim\rho_{i}$. Recall the characteristic flow $\Lambda(t)=\Lambda_{t}$ in (4.4) and choose $T^{(k)}\sim\eta_{*}^{(k-1)}/\rho_{*}=n^{-a_{k-1}+c_{*}}$. From Lemma 4.1, there exist initial conditions $|z_{i}(0)|=|z_{i}^{(k)}|+O(T^{(k)})$ and $|\eta_{i}(0)|\gtrsim\rho_{*}T^{(k)}=\eta_{*}^{(k-1)}=n^{-a_{k-1}}$ such that $\eta_{i}(T^{(k)})=\eta^{(k)}_{i}$ and $z_{i}(T^{(k)})=z^{(k)}_{i}$. The result from the previous step in (5.22) ensures that the assumption (4.18) of Proposition 4.4 (with $b$’ playing the role of $b$) is satisfied for any $|\eta^{(k-1)}_{i}|\gtrsim\eta_{*}^{(k-1)}=n^{-a_{k-1}}$. Thus Proposition 4.4 implies that for $t=T^{(k)}\sim n^{-a_{k-1}+c_{*}}$,
@@
We first apply Itô’s formula to $|R^{xy}_{t}|^{2m}=(R^{xy}_{t})^{m}(\overline{R^{xy}_{t})})^{m}$ and perform the cumulant expansions (see e.g., [56, Lemma 7.1]) as below. Without loss of generality, we may assume in the following that all the cumulants of the normalized entries of $\sqrt{n}X$ are equal to one. We thus obtain
@@
$\mathcal{T}^{(s_{1},\cdots,s_{l})}_{k}:=\frac{1}{n^{\frac{k}{2}}}\sum_{a=1}^{n}\sum_{B=n+1}^{2n}{\mathbf{E}}\Big[\big(\partial^{s_{1}}R^{xy}_{t}\big)\cdots\big(\partial^{s_{l}}R^{xy}_{t}\big)\big(R^{xy}_{t}\big)^{2m-l}\Big],\qquad s_{i}\geq 1,$ (5.27)
where $\partial^{s_{i}}$ denotes the $s_{i}$-th directional derivative either in the direction $h_{aB}$ or in $h_{aB}$, with $(h_{\mathfrak{u}\mathfrak{v}})_{\mathfrak{u},\mathfrak{v}\in[2n]}:=H^{z}_{t}$ given by (5.13). Note that for $a,b\in[n]$ and $A,B\in[n+1,2n]$, the corresponding entries $\{h_{ab}\}$ and $\{h_{AB}\}$ all equal to zero. For notational brevity, we will not distinguish between $\partial h_{aB}$ and $\partial h_{Ba}$ and use the notation $\partial$ to indicate either $\partial/\partial h_{aB}$ or $\partial/\partial h_{Ba}$. From now on, with a slight abuse of notation $R^{xy}_{t}$ indicates either $R^{xy}_{t}\in{\mathbb{C}}$ itself or its complex conjugate $\overline{R^{xy}_{t})}$. Here we stopped the cumulant expansions in (5.1) at a sufficiently high power $K_{0}=O(m)$, and the last error term $O_{\prec}(n^{-100m})$ was estimated using the moment condition in (2.1) and that $\|G\|\leq|\eta|^{-1}$. This argument is quite standard and already used in many previous works, e.g., below [31, Eq. (7.6)], so we omit the details.
@@
with $b^{\prime}=\min\big\{b+\theta,~1\big\}$, $0<\theta\leq\epsilon/10$, using that $\gamma\lesssim\rho^{-2}$ from (3.23) and $\gamma\gtrsim\eta/\rho$.
@@
$|(*)|\prec$ $\frac{\rho^{3/2}}{n\sqrt{\eta}}{\mathbf{E}}\Big[\sqrt{\sum_{B}\big|(G_{1})_{{\bm{x}}{\bm{e}}_{B}}\big|^{2}\sum_{B}\big|(G_{1}AG_{2})_{{\bm{e}}_{B}{\bm{y}}}\big|^{4}}\big|R^{xy}_{t}\big|^{2m-2}\Big]$